# Leveraging Large Language Models (LLMs) to Empower Training-Free Dataset Condensation for Content-Based Recommendation

### v3

@@ Abstract
Modern techniques in Content-based Recommendation (CBR) leverage item content information to provide personalized services to users, but suffer from resource-intensive training on large datasets. To address this issue, we explore the dataset condensation for textual CBR in this paper. The goal of dataset condensation is to synthesize a small yet informative dataset, upon which models can achieve performance comparable to those trained on large datasets. While existing condensation approaches are tailored to classification tasks for continuous data like images or embeddings, direct application of them to CBR has limitations: (1) those methods fail to synthesize discrete texts and (2) they cannot handle the interactions involving both users and items while concurrently preserving users’ preference information during the condensation process.
To bridge this gap, we investigate efficient dataset condensation for content-based recommendation. Inspired by the remarkable abilities of large language models (LLMs) in text comprehension and generation, we leverage LLMs to empower the generation of textual content during condensation. To handle the interaction data involving both users and items, we devise a dual-level condensation method: content-level and user-level. At content-level, we utilize LLMs to condense all contents of an item into a new informative title. At user-level, we design a clustering-based synthesis module, where we first utilize LLMs to extract user interests. Then, the user interests and user embeddings are incorporated to condense users and generate interactions for condensed users. Notably, the condensation paradigm of this method is forward and free from iterative optimization on the synthesized dataset, which is distinguished from previous approaches following the paradigm of nested bi-level optimization. Extensive empirical findings from our study, conducted on three authentic datasets, substantiate the efficacy of the proposed method. Particularly, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND).
Index Terms:
Dataset condensation, Deep Learning, Recommendation
§§footnotetext: Equal contribution
I Introduction
Figure 1: Comparison on the paradigm of dataset synthesis between our TF-DCon and previously proposed methods in other domains.
@@ I Introduction
Content-based recommenders [1, 2, 3] have made strides in mitigating the information overload dilemma, delivering items with relevant contents (i.e., news, articles, or movies) to users. Existing advanced CBR models [4, 5, 6] are trained on large-scale datasets encompassing millions of users and items. Handling these datasets significantly strains computational resources during model training. Furthermore, the recurrent retraining of recommendation models, especially for periodic updates in real-world applications, exponentially elevates costs to unsustainable levels.
The intensifying exploration of dataset condensation, also called dataset distillation, sheld light in addressing aforementioned issues. The goal of dataset condensation is to synthesize a small yet informative dataset trained on which the model can achieve comparable performance to that of a model trained on the original dataset [7, 8, 9, 10, 11, 12]. In particular, one of the most representative methods, DC [11] formulates the condensation on image data as matching gradients of network parameters between synthetic and original data. Doscond [10] continualizes the graph structure and optimizes the synthetic graph data via a one-step gradient matching mechanism. TDD [13] proposes to distill the labels of samples in a soft manner, which could be widely applied in classification tasks and conducts condensation on text embeddings to handle the discrete textual data. In spite of the promising success, dataset condensation on CBR is still unexplored.
To bridge this gap, we investigate how to effectively condense the dataset for a textual content-based recommendation. In the context of CBR, the existing condensation approaches devised in other domains [11, 8, 10, 12] are limited by following challenges:
  • How to generate discrete textual data? Current condensation methods are designed for continuous data (i.e., images and text embeddings), following a formulation of nested bi-level optimization. Under this formulation, the data is synthesized via an outer gradient, which cannot be utilized to generate discrete textual data. Therefore, it is vital for us to find a solution to generate the discrete text.
  • How to handle the interactions between users and items? In the tasks of recommendation, the goal is to utilize the information of users and items as well as the interactions between them to infer the preferences of users. However, previously proposed methods mostly involve classification tasks, which fail to handle the interaction data and condense the preference information within it [14, 15, 16].
While a recent work [14] aims at condensing collaborative filtering dataset, it is only suitable for NTK-based (Neural Tangent Kernel) architecture and fails to process textual data. Besides, CGM [16] follows the nested bi-level formulation to condense categorical data for CTR predictions and DConRec [15] designs an optimization-based method for ID-based recommendation, both of which, however, cannot generate texts as well.
@@
Large Language Models (LLMs) show strong capacity to distill textual information [17, 18, 19], and its expertise across various fields such as medicine [20, 21], recommendation [22, 23, 24], and graph [25]. The possession of versatile nature and extensive world knowledge enables LLMs emergent abilities, such as strong text comprehension and language generation capability [17, 18]. Given its strong text comprehension and language generation capability [17, 18], we are the first to propose utilizing LLMs for processing textual content in condensation for CBR, to address the aforementioned challenges.
@@
Based on previous discussion, we propose a LLMs-empowered Training-Free Dataset Condensation method for content-based recommendation, abbreviated as TF-DCon. Specifically, to handle the interactions and preserve the preference information in condensed dataset [14], TF-DCon is devised as a two-level solution: content-level and user-level. At content-level, we curate prompts to instruct LLMs to condense all information and contents of an item into an informative title. In user-level, we propose a clustering-based synthesis module to simultaneously generate fake users and their corresponding historical interactions, based on user interests extracted by LLMs and user embeddings. Consequently, there are two advantages of our proposed method: (1) TF-DCon omits the conventional paradigm of formulating condensation as a bi-level optimization problem [11, 7, 8]. Instead, we formulate condensation as a forward process, which is free from iteratively training datasets (as shown in Figure 1); (2) the condensed datasets generated by our method are still textual data instead of a set of text embeddings, which makes it more generalizable and flexible to train different architectures of models for CBR. Our contributions could be summarized as follows:
  • To the best of our knowledge, we are the first to investigate the problem of condensing datasets for textual CBR, where we devise a forward condensing paradigm which is different from the widely adopted bi-level optimization formulation.
  • We first propose a LLMs-empowered dataset condensation method. Specifically, we curate prompts to instruct LLMs to condense the contents of items and devise a clustering-based synthesis module to generate fake users and their historical interactions.
@@
  • Extensive empirical findings from our study, conducted on three authentic datasets, substantiate the efficacy of the proposed method. Particularly, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND). The model training on the condensed datasets is significantly faster (i.e., 5×5$\times$5 × speedup).
@@
The rest of the paper is structured as follows. Section V reviews the related works for content-based recommendation and dataset condensation. Section II provides background for relevant preliminaries. Section III elaborates on the details of the proposed condensation method for CBR. Section IV presents the experimental results and analysis. The conclusion is obtained in section VI.
II Preliminaries
@@
Figure 2: The proposed method in a nutshell. The LLMs-empowered content-level condensation and user interest extraction are instanced in Figure 3.
II-A Content-based Recommendation
@@
Before elaborating on our proposed method, let us introduce some basic knowledge and a formal definition of content-based recommendation. CBR aims to infer the preference of users based on historical interactions and then recommend the contents with potential interest. In the context of CBR, the dataset $\mathcal{D}$ mainly consists of item set $\mathcal{N}$, user set $\mathcal{U}$, and click history set $\mathcal{H}$. Each item $n\in\mathcal{N}$ contains various contents, such as title, abstract, and category. Each user $u\in\mathcal{U}$ is associated with a clicking history of items $h^{(u)}\subset\mathcal{N}$. We denote the set of history $h^{(u)}$ by $\mathcal{H}$. Let $\mathcal{C}$ denote the set of clicks, where each click $c\in\mathcal{C}$ is defined as a tuple $(u,n,y)$ indicating whether user $u$ clicked on item $n$ with label $y\in\{0,1\}$. The task of CBR is to infer the preference of users on given candidate items based on their historical interactions.
II-B Dataset Condensation
@@
III Method
III-A Overview of the Proposed Method
In this section, we detail the proposed method. Due to the inherent discreteness of textual data, the conventional gradient-based dataset condensation approaches, which update the synthetic data via the outer gradient backpropagation in the bi-level formulation, is invalid for condensing textual contents. Fortunately, the emergent abilities of LLMs [17, 18], such as strong text comprehension and language generation capabilities, shed light on this issue. Hence, we utilize LLM to assist the condensation. To correspondingly address the challenges, we discuss the approaches in section III-B and section III-C. In section I, we will discuss comparison on the time complexity of different condensation methods.
As shown in Figure 2, the proposed method condenses the content-based recommendation datasets from two levels: content-level and user-level. In content-level condensation, we aim to alleviate the textual data load for each item by condensing its content (Section III-B). In user-level condensation, we aim to minimize historical redundancies by pinpointing users’ primary interests and merging users with similar preferences (Section III-C). Both content compression and user synthesis leverage the capabilities of advanced LLMs, which we’ll delve into in the subsequent sections.
III-B Content-Level Condensation
In the content-level condensation, we aim to condense all the information of each item into a succinct yet informative title. Recent studies [26, 27] reveal the exceptional power of large language models in processing textual contents. Therefore, we propose to utilize LLM to condense the contents. Specifically, we design the guiding prompt in the following format:
@@ III-B Content-Level Condensation
After facilitating LLM’ comprehension on the input, the information of items will be fed into LLM to generate the condensed title. Formally, the process can be described as follows:
$n_{s}=LLM(n),$ (2)
@@
III-C User-Level Condensation

### v4

@@ Abstract
Modern techniques in Content-based Recommendation (CBR) leverage item content information to provide personalized services, effectively alleviating information overload. However, these methods suffer from resource-intensive training on large datasets. To address this issue, we explore the dataset condensation for textual CBR in this paper. Dataset condensation aims to synthesize a compact yet informative dataset that enables models to achieve performance comparable to those trained on full datasets. Directly applying existing approaches to CBR has limitations: (1) difficulties in synthesizing discrete texts and (2) the inability to retain preference information between users and items during condensation. To bridge this gap, we propose TF_DCon, an efficient dataset condensation method for CBR. Leveraging ChatGPT, TF_DCon enables condensation on discret texts and preserves preference via a clustering-based synthesis module to condense users. Unlike previous methods, TF_DCon’s forward pipeline eliminates the iterative updates on the condensed data, greatly enhancing the efficiency. Extensive experiments conducted on three real-world datasets demonstrate TF_DCon’s effectiveness. Notably, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on dataset MIND).
keywords:
[…]
1 Introduction
@@
Content-based recommenders (Wu et al., 2020, 2019a, 2021) have made strides in mitigating the information overload dilemma, delivering items with relevant content (i.e., news, articles, or movies) to users. Existing advanced content-based recommendation (CBR) models (Wang et al., 2018a; Wu et al., 2019b) are trained on large-scale datasets encompassing millions of users and items. Handling these large datasets significantly strains computational resources during model training. Furthermore, the recurrent retraining of recommendation models, especially for periodic updates in real-world applications, exponentially elevates costs to unsustainable levels.
The intensifying exploration of dataset condensation, also called dataset distillation, shed light in addressing the aforementioned issues. The goal of dataset condensation is to synthesize a small yet informative dataset, trained on which the model can achieve comparable performance to that of a model trained on the original dataset (Wang et al., 2018b; Jin et al., 2022b; Zheng et al., 2023; Jin et al., 2022a; Zhao et al., 2021; Han et al., 2022). The most prevalent paradigm (Zhao et al., 2021; Jin et al., 2022a; Sucholutsky and Schonlau, 2021) is to formulate the condensation as bi-level optimization problem and iteratively update the condensed data by matching gradients of network parameters between synthetic and original data. Such methods have achieved success on condensing continuous data, such as images and node features. However, condensing datasets for CBR is still unexplored.
Figure 1: Comparison on the pipelines of dataset synthesis between our TF-DCon and previously proposed methods in other domains.
To bridge this gap, we investigate how to effectively condense the dataset for the textual content based recommendation. Existing condensation approaches devised in other domains (Zhao et al., 2021; Jin et al., 2022b, a; Han et al., 2022) face two main challenges in the context of CBR: (1) How to generate discrete textual data? Current condensation methods are designed for continuous data (i.e., images and text embeddings), following a formulation of nested bi-level optimization. Under this formulation, the data is synthesized via the outer gradient, which cannot be utilized to generate discrete textual data. Therefore, a solution for text synthesis is essential. (2) How to preserve the preference information between users and items? In recommendation tasks, user and item data, along with their interactions, are critical for inferring user preferences. However, previously proposed methods mostly involve classification tasks, limiting their ability to handle interaction data and retain preference information.
@@
Large Language Models (LLMs), such as ChatGPT developed by OpenAI, have shown a strong capacity to distill textual information (Brown et al., 2020; Chowdhery et al., 2022; Li et al., 2023a), and its general expertise across various fields such as medicine (Antaki et al., 2023; Benoit, 2023), recommendation (Fan et al., 2023; Dai et al., 2023b), and law (Cui et al., 2023a; Choi et al., 2023). With their versatility and extensive world knowledge, LLMs exhibit emergent abilities such as robust text comprehension and language generation (Brown et al., 2020; Chowdhery et al., 2022). Leveraging these capabilities (Brown et al., 2020; Chowdhery et al., 2022), we propose to leverage ChatGPT for dataset condensation, processing textual content to address one of the aforementioned challenges.
@@
While LLMs are highly effective at processing text, they lack specific domain knowledge for recommendation scenarios and the ability to capture personalized user preferences. Naïvely applying LLMs for recommendation data condensation risks losing both this domain knowledge and essential preference information. To this end, we propose a ChatGPT-powered Training-Free Dataset Condensation method for content-based recommendation, abbreviated as TF-DCon. TF-DCon is devised in a two-level manner: content-level and user-level. At content-level, we curate a prompt-evolution module to optimize prompts, enabling ChatGPT to adapt to the specific recommendation domain and condense each item’s information into an informative title. At user-level, to capture the preference information of users, we propose a clustering-based synthesis module to simultaneously generate fake users and their corresponding historical interactions, based on user interests extracted by ChatGPT and user embeddings. Consequently, our approach offers several advantages: (1) TF-DCon omits the conventional paradigm of formulating condensation as a bi-level optimization problem (Zhao et al., 2021; Wang et al., 2018b; Jin et al., 2022b). Instead, we formulate condensation as a forward process, eliminating iteratively training on datasets and greatly enhancing the condensation efficiency (see Figure 1); (2) TF-DCon can generate text, which makes the condensed dataset more generalizable and flexible to train different architectures of models for CBR. (3) TF-DCon seamlessly embeds user preferences into the condensed data through the clustering-based synthesis module. Our contributions are summarized as follows:
  • Objective. To the best of our knowledge, our study presents the first exploration of dataset condensation for textual CBR. We introduce a forward condensation paradigm distinct from traditional bi-level optimization and aim to effectively condense the information of items and user preference.
  • Methodology. We propose a ChatGPT-driven dataset condensation method featuring a prompt-evolution module for item content condensation and a clustering-based synthesis module to generate synthetic users and interactions, thereby preserving both item and preference information effectively.
@@
  • Experiment. Extensive experimental results demonstrate the efficacy of the proposed method. In particular, we are able to approximate up to $97\%$ of the original performance while reducing the dataset size by $95\%$ (i.e., on the dataset MIND). The model training on the condensed datasets is significantly faster (i.e., 5×5$\times$5 × speedup).
@@
The rest of the paper is structured as follows. Section 2 reviews the related works for content-based recommendation and dataset condensation. Section 3 provides background for relevant preliminaries. Section 4 elaborates on the details of the proposed condensation method for CBR. Section 2 presents the experimental results and analysis. Section 6 discusses the limitations of our proposed method. The conclusion is obtained in section 7.
2 Related Work
2.1 Content-Based Recommendation
Content-based recommendation (CBR) involves a wide range of information in different formats: text (Li et al., 2023b; Wu et al., 2022), image (Wang et al., 2024a; Zhang et al., 2024), video (Wang et al., 2024b; Xiao et al., 2024), social networks (Gan et al., 2024; Liu et al., 2024), etc. In this paper, we mainly focus on textual contents, i.e., news, books, and movie reviews.
In recent years, textual CBR has garnered substantial attention and has been the subject of comprehensive investigation within both industrial and academic domains. Models based on neural networks (Wang et al., 2018a; Wu et al., 2019b) have been proposed to learn the representations of users and items via CNN (Krizhevsky et al., 2012) or attention mechanism (Vaswani et al., 2017). Thereafter, the advance of pretrained language models like BERT (Devlin et al., 2019) exhibited promising performance on textual content recommendation (Li et al., 2022; Zhang et al., 2021). As those methodologies depend on training with extensive-scale datasets, their exceptional performance comes at a high computational cost, which pose a challenge for sustainable model training and further scalable application. The recent emergence of investigations into dataset condensation and dataset distillation presents a promising avenue for tackling this concern (Wang et al., 2018b; Zhao et al., 2021). The inherent concept underlying dataset condensation and dataset distillation revolves around the synthesis of a compact dataset for model training, with the objective of enabling models trained on this condensed dataset to achieve performance levels comparable to those trained on the complete dataset. Nevertheless, no extant research has delved into the realm of condensation specifically within the context of CBR. To bridge this critical gap, we introduce an innovative and efficacious framework tailored to the task of dataset condensation for CBR.
2.2 Dataset Condensation
It is notorious that training neural networks on large datasets is prohibitively expensive. To mitigate this issue, plenty of works on dataset distillation/dataset condensation have been proposed to synthesize small datasets to replace the large ones, such as GCond (Jin et al., 2022b), DC (Zhao et al., 2021) and SFGC (Zheng et al., 2023). Those methods always formulate the condensation process as a bi-level optimization problem. The condensation methods have been explored on images (Zhao et al., 2021), graph (Jin et al., 2022b) and text embeddings (Han et al., 2022). Various methods have been proposed to improve the quality of condensation and optimization efficiency. For instance, DC (Zhao et al., 2021) proposed gradient matching, achieving a comparable performance under a high compression ratio, and DosCond (Jin et al., 2022a) proposed one-step update strategy, which significantly improves the optimization efficiency. Condensation on textual data is conducted based on the text embeddings or soft labels (Maekawa et al., 2023; Li and Li, 2021; Sahni and Patel, 2023; Sucholutsky and Schonlau, 2021; Han et al., 2022; Wang et al., 2023b). Further exploration extends condensation methods to image-text retrival (Wu et al., 2023b). Recently, a work (Sachdeva et al., 2022) aims at condensing collaborative filtering dataset but it is only suitable for NTK-based (Neural Tangent Kernel) architecture and fails to process textual data. Besides, CGM (Wang et al., 2023a) follows the nested bi-level formulation to condense categorical data for CTR predictions and DConRec (Wu et al., 2023a) designs an optimization-based method for ID-based recommendation, both of which cannot generate texts as well.
Despite the extensive investigation into dataset condensation, little attention has been paid to the domain of CBR. To address this disparity, we design a condensation method for the CBR dataset. Unlike earlier methodologies, our dataset condensation approach is devoid of training requirements, leading to a substantial reduction in condensation expenses. Furthermore, our condensation procedure centers on textual content, thereby diverging from those text condensation methods that hinge on text embeddings.
3 Preliminaries
@@
Figure 2: The Proposed Method in a Nutshell. The outline of this method is presented in Algorithm 1 and the details of the prompt’s evolution are outlined in Algorithm 2. “N. G.” denotes “Next Generated Prompt Candidates”. Given a prompt, the content-level condensation and user interest extraction are instanced in Figure 3.
3.1 Content-based Recommendation
@@
Before elaborating on our proposed method, let us introduce some basic knowledge and a formal definition of content-based recommendation. CBR aims to infer the preference of users based on historical interactions and then recommend the contents with potential interest. In the context of CBR, the dataset $\mathcal{D}$ mainly consists of item set $\mathcal{N}$, user set $\mathcal{U}$, and click history set $\mathcal{H}$. Each item $n\in\mathcal{N}$ contains various contents, such as title, abstract, and category. Each user $u\in\mathcal{U}$ is associated with a clicking history of items $h^{(u)}\subset\mathcal{N}$. We denote the set of history $h^{(u)}$ by $\mathcal{H}$. Let $\mathcal{C}$ denote the set of clicks, where each click $c\in\mathcal{C}$ is defined as a tuple $(u,n)$ indicating that user $u$ has clicked on item $n$. The task of CBR is to infer the preference of users on given candidate items based on their historical interactions.
3.2 Dataset Condensation
@@
  Require: dataset $\mathcal{D}=(\mathcal{H},\mathcal{N})$ where $\mathcal{H}$ is the set of users’ historical interactions and $\mathcal{N}$ is the set of items, number of synthesized users in the condensed dataset $K$. 1:  Utilize ChatGPT to extract user interests: $\mathcal{I}\leftarrow ChatGPT(\mathcal{H},\mathcal{N})$. 2:  Encode user interests and user histories:$H\leftarrow UserEncoder(\mathcal{D})$, $H^{\prime}\leftarrow TextEncoder(\mathcal{I})$. 3:  Clustering to generate $K$ synthetic users and their interacted item lists: $\mathcal{H}^{con}\leftarrow Clustering(K,H,H^{\prime})$. // Section 4.2. 4:  Utilize ChatGPT to condense the content of items: $\mathcal{N}^{con}\leftarrow ChatGPT(\mathcal{N})$. // Section 4.1.   Return Condensed dataset $\mathcal{D}^{con}=(\mathcal{H}^{cond},\mathcal{N}^{cond})$.
Algorithm 1 TF-DCon
4 Method
In this section, we detail the proposed method (TF-DCon) and the overview is shown in Figure 2. Content-level condensation reduces the textual data load for each item (Section 4.1), while user-level condensation synthesizes fake users and interactions, capturing primary interests and merging users with similar preferences (Section 4.2). Section 1 compares the computational complexity of different condensation methods.
4.1 Content-Level Condensation
4.1.1 Content Condensation
In the content-level condensation, we aim to condense all the information of each item into a succinct yet informative title. Recent studies (Dai et al., 2023a; Ubani et al., 2023) reveal the exceptional power of large language models in processing textual contents. Therefore, we propose to utilize ChatGPT to condense the contents. Specifically, we design the guiding prompt in the following format:
@@ 4.1.1 Content Condensation
After facilitating ChatGPT’s comprehension on the input, the information of items will be fed into ChatGPT to generate the condensed title. Formally, the process can be described as follows:
$n_{s}=ChatGPT(n),$ (2)
@@
4.1.2 Prompt Evolution
To efficiently guide ChatGPT to adapt to the recommendation scenario and perform thorough content condensation for each item, we propose a curated prompt evolution method, outlined in Algorithm 2. Given an initial prompt, EvoPro can guide the prompts to evolve for pre-defined times. We name the initial prompt as 0-th generation prompt, abbreviated as $gen_{0}$-prompt. During $i$-th iteration of evolving, EvoPro will first instruct ChatGPT to generate $N$ next-generation prompts $\{prompt_{n}\}$ based on $gen_{i-1}$-prompt. Then, these prompts will be utilized to instruct ChatGPT to condense item contents for TF-DCon. The similarity scores between the embeddings of condensed contents and the original contents will be assigned to each prompt in $\{prompt_{n}\}$, the calculation of which is outlined in Algorithm 3. The prompt with the highest score will be selected as $gen_{i}$-prompt for $i$-th generation. When the evolution finished, the prompt with the hightest score in the latest generation will be selected as the prompt for content condensation.
The rationale behind the score for prompt selection is that the condensed title sharing the largest similarity with the original contents preserves the most information.
  Require: initial prompt $gen_{0}$-prompt, item contents $\mathcal{C}=\{content_{i}\}$, evolving times $E$, number of prompts each generation $N$. 1:  Initial prompts set $\mathcal{P}\leftarrow${} and counter $e\leftarrow 0$ 2:  while $e<E$ do 3:     $e\leftarrow e+1$. 4:     Use ChatGPT to generate $gen_{e-1}$-prompt’s $N$ next generation prompts $\{prompt_{n}\}$. 5:     Initialize the score set $\mathcal{S}\leftarrow\{\}$. 6:     for all $prompt$ in $\{prompt_{n}\}$ do 7:        Calculate score for $prompt$: $s\leftarrow CalScore(prompt,\mathcal{C})$.// Algorithm 3. 8:        $\mathcal{S}\leftarrow\mathcal{S}\cup\{s\}$ 9:     end for 10:     Select the prompt with the highest score in $\mathcal{S}$ and set it as $gen_{e}$-prompt. 11:     $\mathcal{P}\leftarrow\mathcal{P}\cup\{gen_{e}\text{-}prompt\}$. 12:  end while   Return evolved prompts $\mathcal{P}$.
Algorithm 2 EvoPro
  Require: $prompt$, item contents $\mathcal{C}$. 1:  $s\leftarrow 0$. 2:  for all $c_{i}$ in $\mathcal{C}$ do 3:     Condense item content: $con_{i}\leftarrow LLM(prompt,c_{i})$. 4:     $h_{i}\leftarrow TextEncoder(c_{i}),h^{\prime}_{i}\leftarrow TextEncoder(con_{i})$. 5:     Calculate the similarity score: $s\leftarrow s+Sim(h_{i},h^{\prime}_{i})$. 6:  end for   Return $s$
Algorithm 3 CalScore
4.2 User-Level Condensation